OpenStudy (anonymous):

P = one-sixth (r + 5t) for r

OpenStudy (anonymous):

Just make r alone at the right hand side

OpenStudy (anonymous):

wait so what do i have to do after that

OpenStudy (anonymous):

that will be your answer..... if u do that... for example: P=2+r for r P=2+r P-2=r r=P-2

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so what would my outcome be

OpenStudy (anonymous):

P=1/6 (r+5t) 6P=r+5t

OpenStudy (anonymous):

Now can u?

OpenStudy (anonymous):

yes i got the answer r=one/sixth+5t

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

really

OpenStudy (anonymous):

6p=r+5t r=6p-5t

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

oh man ok i get it now